THE MEASURE OF ONE ANGLE OF A TRIANGLE IS 10 DEGREES MORE THAN THE MEASURE OF THE SMALLEST ANGLE. THE MEASURE OF THE THIRD ANGLE IS 50 DEGREES MORE THAN THE MEASURE OF THE SMALLEST ANGLE. WHAT ARE THE MEASURES OF THE ANGLES OF THE TRIANGLE?
Answer by checkley77(12844)

x, x+10 & x+50 are the three angles.
x+x+10+x+50=180
3x=180-10-50
3x=120
x=120/3
x=40 degrees for the smaller angle.
40+10=50 degrees for the middle angle.
40+50=90 degrees for the largest angle.
Proof:
40+50+90=180
180=180